A stone is thrown horizontally from the top of a vertical cliff and lands 4.0s later at a distance
12.0m from the base of the cliff. Ignore air resistance.
a) Calculate the horizontal speed of the stone.
b) Calculate the height of the cliff.

Answer by Alan3354(69443)

A stone is thrown horizontally from the top of a vertical cliff and lands 4.0s later at a distance 12.0m from the base of the cliff. Ignore air resistance.
a) Calculate the horizontal speed of the stone.
12/4 = 3 m/sec
-----
b) Calculate the height of the cliff.
h = at^2/2
Using 9.8 m/sec/sec for gravity:
h = 9.8*4^2/2 meters